A corporate bond with a face value of $1,000 matures in 4 years and has a coupon rate of 6.25 percent. The current price of the bond is $932 and interest is paid semiannually. If inflation averaged 3.26 percent, what was the real rate of return?
Answer #1

So, let us first calculate the yield to maturity of the bond,

FV = $1000

N = 4 * 2 = 8 YEARS ( Semi-annual compounding)

PMT = 6.25% / 2 * $1000

= $31.25

PV = ($932)

I/ Y = 4.1409

So, the nominal rate of interest is = 8.2817 ( 4.1409* 2)

So,as per the Fisher formula,

(1 + nominal rate of interest) = (1 + real rate of interest) ( 1 + inflation )

(1 + 0.0828) = (1 + real rate of interest) (1 + 0.0326)

Real rate of interest = 4.86%
It can also be calculated as , 8.28% - 3.26%

So, the real rate of interest is 5.02%.